Problem Statement 3 Every spring semester, the School of Business coordinates a luncheon with local business leaders for graduating seniors, their families, and friends. Corporate sponsorship pays for the lunches of each of the seniors, but students have to purchase tickets to cover the cost of lunches served to guests they bring with them. The following histogram represents the attendance at the senior luncheon, where X is the number of guests each graduating senior invited to the luncheon and f is the number of graduating seniors in each category.

preview

Referring to the histogram from Problem Statement 3, how many graduating seniors attended the luncheon?

Referring to the histogram from Problem Statement 3, if all the tickets purchased were used, how many guests attended the luncheon?
Answer #1

A bs olute Frequency Histo g ram 200 150 100 152 85 5017 Number of Gu e sts

(a) The number of graduating seniors is the sum of all the frequencies, f.

= 17+152+85+18+3=275.

(b) The total number of guests is summation of product of X and f

= 17*0 + 152*1 + 85*2 + 18*3 + 3*4

= 388